OpenStudy (anonymous):

derivative of these: F(x)=-1/4 cos2x G(x)=-1/2 cos^2 x

OpenStudy (raden):

actually, the chain rule will works

OpenStudy (raden):

for F(x)=-1/4 cos2x the derivative oc 2x = 2 the derivative of cos(...) = -sin(...) so, f ' (x) = -1/4 (2-sin(2x)) = - 1/2 + (1/4) sin(2x)

OpenStudy (raden):

* of

OpenStudy (anonymous):

right ............... im trying it now

OpenStudy (anonymous):

the chain rule made me see the idea clearly ... thanks